# Data-Driven Goal Recognition Design for General Behavioral Agents

### v1

@@ 4.2 Baselines
    Pruned-Reduce (Keren et al., 2014): This baseline is specifically designed for settings where modifications are limited to action removal such as blocking a cell in grid world. It speeds up the exhaustive search and retains the optimal property. However, its scalability is still limited.
@@ 4.3 Models of Agent Behavior
    Parameterized suboptimal agent behavior. We also consider an generalized behavior model that is captured by a time-variant discounting function $d(t)$, and the agent’s objective is to optimize a time discounted reward with the discounting factor for $t$ steps in the future being $d(t)$. The standard optimal agent model corresponds to a fixed discounting factor $\gamma\in(0,1]$ with $d(t)=\gamma^{t}$. We can also represent an agent with present bias (O’Donoghue and Rabin, 1999) by adopting a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$, where $k>0$.
@@ 5 Experiments
We have conducted extensive simulations to examine the performance of our approach. Moreover, we have conducted real-world human-subject experiments to assess whether our approach is effective with human decision-makers.
@@ 5.1.1 Standard setting
We first focus on settings within the grid world domain, where modifications are limited to blocking cells and agent behavior is assumed to be optimal. This scenario forms the standard setting for the majority of goal recognition design studies, as highlighted in Table 1 of the survey by Keren et al. (2021). The objective of this set of simulations is to enable comparisons with state-of-the-art methods.
In our experiments, the initial environment is generated randomly: we first randomly select the number of blocked cells from the range $[0,12]$, followed by the random allocation of the blocked cells’ positions. We also randomly determine the starting grid and two goal grids. Environments where the goals are not reachable from the starting grid are filtered out. Following this procedure, we randomly generate 200 environments and compare the average performance of our approach with that of baseline methods.
@@
Figure 3: wcd reduction in a grid world when only blocking modifications are allowed. Exhaustive search and Pruned-Reduce are not included in the larger environments of (b) because they cannot complete the computation within an hour for a single environment.
@@ 5.1.2 Settings with flexible budget constraints
(a) Shared budget.
(b) Individual budget.
Figure 4: wcd reduction in settings with two types of modifications.
@@ 5.1.3 Complex domain and suboptimal agent behavior
Figure 6: Grid world of size $6\times 6$. Incorporating suboptimal agent behavior.
@@ 5.2.1 Experiment 1: Collection of human behavioral data
Figure 7: The interface of our human subject experiment.
Table 1: The prediction accuracy for human behavior assuming optimal behavior and using data-driven model in Experiment 1.
Training Accuracy Validation Accuracy Testing Accuracy
Assuming Optimal Behavior 0.7266 0.6964 0.7131
Data-Driven Model 0.9189 0.8136 0.8422
Learning models of human behavior.
The collected human data were divided into training, validation, and testing sets: the training set included data from 160 workers, comprising approximately 70,000 instances of user decisions, while the validation and testing sets each contained data from 20 workers, amounting to around 8,800 instances of user decisions each. We employed a 4-layer Multilayer Perceptron (MLP) model for training the model. The input to the model is the current environment layout, and its output predicts the next human action. We fine-tuned the hyperparameters, including learning rate, hidden layer size, and L2 penalty, based on the performance on the validation dataset.
We compared the performance of our learned model against a model that assumes optimal agent behavior, which is defined as taking the shortest path to the goal. The training, validation, and test accuracies of both models are presented in Table 1. These results clearly indicate that human behavior deviates significantly from optimality. This deviation underscores the importance of incorporating a realistic model of human behavior in goal recognition design, particularly when humans are the decision-makers.
@@ 5.2.2 Experiment 2: Evaluating goal recognition design
    Initial: No updates to the environment.
@@
We recruited 200 workers from Amazon Mechanical Turk. Each worker was randomly assigned to one of the four treatments specified above, with the distinction between treatments being the environments presented to them. Workers were randomly assigned a goal for each environment and tasked with navigating to reach it. After collecting their series of actions aimed at reaching the goal, we analyzed this behavioral data to assess the effectiveness of goal recognition design for each approach.
Comparing average overlapping path.
To evaluate the effectiveness of each goal recognition design approach, we first measure the average overlapping path towards each of the two goals. Specifically, each recruited worker is exposed to all 30 environments in their assigned treatment and is instructed to reach one of two randomly selected goals. For each treatment, we calculate the number of overlapping actions for every pair of workers assigned to different goals within each environment. We then compute the average of these numbers, yielding the average overlapping path as our performance measure. This metric reflects the difficulty of inferring agent goals for the envioronment and serves as an empirical proxy for wcd. With this metric, we vary the initial $k$ portion of actions taken by the workers in our analysis and observe how the average overlapping path changes with the amount of observed actions, denoted by $k$. As depicted in Figure 10(c), our approach, when integrated with data-driven models, leads to environments that facilitate the easier identification of the humans’ intended goal.
Comparing the accuracy of goal inference.
We next measure the accuracy of inferring human goals based on partial observations, utilizing an off-the-shelf Bayesian inference algorithm [^3] We assume an initial uniform prior distribution. As the agent takes actions, we update the posterior belief over goals based on the likelihood probability of the data-driven behavior model prediction. . Specifically, for each worker, we reveal the first $k$ portion of the worker’s actions to the inference algorithm to infer the worker’s goal. This enables us to compute the average inference accuracy. The results, as shown in Figure 8(b), demonstrate that our approach leads to environments that are easier for goal recognition.
(a) Averaging overlapping path (the lower the better.)
@@ 6 Conclusion
Effective collaboration between humans and AI hinges on the AI’s ability to infer intentions and goals of humans. In this study, we work on the problem of goal recognition design, aiming to update the decision-making environments to make it easier for the AI agents to perform goal recognition. By addressing the key limitations in the existing literature, notably the computational demand and the assumption of optimal agent behavior, we have developed a data-driven optimization framework that is both efficient and scalable. Moreover, our approach is capable of integrating general models of various agent behavior, moving beyond the traditional assumption of optimality and making it more applicable to real-world scenarios involving human agents.
Future work of our study includes extending our framework to more complex environments, particularly those that are dynamic and partially observable. Given our framework is adaptable to different forms of agent behavior, understanding how different patterns of agent suboptimality impacts the difficulty of goal recognition could lead to useful insights for enabling better human-AI collaboration.
@@ Appendix B Additional Experiment Results
Appendix C Simulation Details
@@ Appendix D Details of Human-Subject Experiment
Figure 12: Human-Subject experiment interface.

### v2

@@ 4.2 Baselines
    Pruned-Reduce [12]: This baseline is specifically designed for settings where modifications are limited to action removal such as blocking a cell in grid world. It speeds up the exhaustive search and retains the optimal property. However, its scalability is still limited.
@@ 4.3 Models of Agent Behavior
    Parameterized suboptimal agent behavior. We also consider an generalized behavior model parameterized by $d(t)$. The agent’s objective is to optimize a time discounted reward with the discounting factor for $t$ steps in the future being $d(t)$. The standard optimal agent model corresponds to a fixed discounting factor $\gamma\in(0,1]$ with $d(t)=\gamma^{t}$. We can also represent an agent with present bias [19] by adopting a hyperbolic discounting factor $d(t)=\frac{1}{1+kt}$, where $k>0$.
@@ 5.1.1 Standard setting
We first focus on settings within the grid world domain, where modifications are limited to blocking cells and agent behavior is assumed to be optimal. This scenario is the standard setting for the majority of goal recognition design studies, as highlighted in Table 1 of the survey by Keren et al. [15]. The objective of this set of simulations is to enable comparisons with state-of-the-art methods.
In our experiments, the initial environment is generated randomly: we first randomly select the number of blocked cells from the range $[0,12]$, followed by randomly allocating the blocked cells. We also randomly determine the starting grid and two goal grids. Environments where the goals are not reachable from the starting grid are filtered out. We randomly generate 200 environments and compare the average performance of our approach with that of baseline methods.
@@
Figure 3: wcd reduction in a grid world when only blocking modifications are allowed. Exhaustive search and Pruned-Reduce are not included in (b) because they take more than an hour to compute for a single environment.
(c) Shared budget
(d) Individual budget
Figure 4: wcd reduction in settings with two types of modifications. We only included greedy baselines as the state-of-the-art baselines, such as Exhaustive search and Pruned-Reduce, are not applicable in these settings.
@@
Figure 6: Grid world ($6\times 6$). Incorporating suboptimal agent behavior.
@@ 5.2.1 Experiment 1 : Collection of human behavioral data
Learning models of human behavior. The collected human data were divided into training, validation, and testing sets: the training set included data from 160 workers, with approximately 70,000 instances of user decisions, while the validation and testing sets each contained data from 20 workers, with around 8,800 instances of user decisions each. We employed a 4-layer Multilayer Perceptron (MLP) model for training the model. The input to the model is the current environment layout, and its output predicts the next human action. We fine-tuned the hyperparameters based on the performance on the validation dataset. We compared the performance of our learned model against a model that assumes optimal agent behavior, which is defined as taking the shortest path to the goal. The training, validation, and test accuracies of both models are presented in Table 1. These results clearly indicate that human behavior deviates significantly from optimality. This deviation underscores the importance of incorporating a realistic model of human behavior in goal recognition design, particularly when humans are the decision-makers.
Table 1: Prediction accuracy of human behavior assuming optimal behavior versus data-driven model.
Assuming Optimal Behavior Using Data-Driven Model
Training accuracy 0.7266 0.9189
Validation accuracy 0.6964 0.8136
Testing accuracy 0.7131 0.8422
@@ 5.2.2 Experiment 2: Evaluating goal recognition design
    Original: No updates to the environment.
@@
We recruited 200 workers from MTurk. Each worker was randomly assigned to one of the four treatments above, with the distinction between treatments being the environments presented to them. Workers were randomly assigned a goal for each environment and tasked with navigating to reach it. We utilize the collected data to evaluate different goal recognition design approaches.
(a) Empirical overlapping actions (the lower the better).
@@
Comparing empirical overlapping actions. To evaluate the effectiveness of each goal recognition design approach, we first measure the empirical overlapping actions towards each of the two goals. Specifically, each recruited worker is exposed to all 30 environments in their assigned treatment and is instructed to reach one of two randomly selected goals. For each treatment, we calculate the number of overlapping actions for every pair of workers assigned to different goals within each environment. Therefore, for each treatment, we obtain a distribution of the number of overlapping actions. This metric reflects the difficulty of inferring agent goals for the environment and serves as an empirical proxy for wcd. Figure 7(a) presents the percentiles of the number of overlapping actions for each treatment. In particular, in the lowest 25%, the environments by our approaches have no overlapping actions, indicating that the environment makes it easy to infer the goal of the agents from observations. Generally, our approach, when integrated with data-driven models, leads to environments with fewer number of overlapping actions and therefore facilitate the easier identification of the humans’ intended goal.
Comparing the accuracy of goal inference. Instead of using wcd or proxy of wcd for evaluations, we next directly measure whether we can indeed infer the human goals based on partial observations of their actions, utilizing an off-the-shelf Bayesian inference algorithm. [^3] We assume an initial uniform prior distribution. As the agent takes actions, we update the posterior belief over goals based on the likelihood probability of the data-driven behavior model prediction. Specifically, for each worker, we reveal the first $k$ portion of the worker’s actions to the inference algorithm to infer the worker’s goal. This enables us to compute the average inference accuracy. The results in Figure 7(b) demonstrate that our approach leads to environments that are easier for goal recognition.
@@ 6 Conclusion
Effective human-AI collaboration hinges on the AI’s ability to infer the goals of humans. In this study, we work on the problem of goal recognition design, updating the decision-making environments to make it easier for the AI agents to perform goal recognition. By addressing the key limitations in the existing literature, notably the computational demand and the assumption of optimal agent behavior, we have developed a data-driven optimization framework that is both efficient and scalable. Through simulations and human-subject experiments, we show that our approach outperforms state-of-the-art approaches in standard settings, applies to settings that existing approaches cannot address, and leads to environments easier for goal recognition with real-world human decision makers.
Societal impact. Our approach is capable of integrating general models of various agent behavior, making it more applicable to real-world scenarios involving human agents. However, with more accurate human models, it also opens up the concerns of privacy and potential mis-use of the techniques for unethical surveillance. While our work does not directly facilitate such misuse, the broader field of goal recognition research could potentially be exploited.
Limitations and future work. While our work has extended the literature in goal recognition design to address suboptimal agent behavior, more flexible budget constraints, and more complicated environments, there are still several limitations. In particular, we have assumed static and full observability of the environment. Extending our framework to deal with more complex environments, particularly those that are dynamic and partially observable would be an interesting and important future work. Moreover, we have assumed that user behavior remains static, so we can learn a human model from their historical behavior. Finally, given our framework is adaptable to different forms of agent behavior, understanding how different patterns of agent suboptimality impacts the difficulty of goal recognition could lead to useful insights for enabling better human-AI collaboration.
Acknowledgements
This work is supported in part by J.P. Morgan Faculty Research Award and a Global Incubator Seed Grant from McDonnell International Scholars Academy.
@@ A.2 Results for Run Time Comparisons
All our experiments are conducted on a cluster of 40 CPU cores (Intel Xeon Gold 6148 CPU @@@@2.40GHz), 1 GPU (NVIDIA Tesla V100 SXM2 32GB), and a maximal memory of 80GB.
@@ Appendix B Additional Experiment Results
B.1 Definition of Budget
In our evaluation, we compare the wcd reduction relative to the budget allocated for modifications. In the grid-world domain, the budget represents the number of changes made, which are limited to two types: blocking or unblocking cells. In contrast, the Overcooked-AI domain allows for a richer set of modifications due to its complexity. Here, modifications involve changing the positions of objects. A valid environment must include all specified objects, as detailed in Section 4.1.2. The budget in this domain is quantified as the total Manhattan distance moved by the objects between the original and modified environments.
Appendix C Details of Simulations
@@ Appendix D Details of Human-Subject Experiment
Note that while each environment has two goals, we only show one goal (the goal of the worker) to the worker in our interface to simplify the presentations. The second goal is shown as a blocked cell in the interface, i.e the worker only navigates to the shown goal.
Figure 11: The interface of our human subject experiment.